a bullet fired vertically up from the ground reaches a height 40 m in its path from the ground and it takes further time 2 seconds to to reach same point during descent.the total time of flight is (g=10 ms-²) .

Grade:11

2 Answers

shubhambhargava
6 years ago
As time taken to reach same point is given I.e 2 sec so at that point velocity must be 10 m/s by 2u/g=t which 20/ 2 =10m/s =u and also height when the velocity is 10 is given which is 40 metre so by v^2-u^2=2gh we got initial velocity 30 m/s.so then the total time of flight is 2u/g that is 6sec.hope you got your answer.
BHARGAVA MADHUKAR
4 years ago
time taken to reach same point is given I.e 2 sec is equal to 2*velocity at that time / g. So     2*v/10 =2.      
V=10m/s
The height reached by the particle i.e 40 m
(10)^2 - u ^2 = 2(-g)40
u^2=900
U=30 m/s
Time of flight = 6s